### v1

# ConfQA: Answer Only If You Are Confident

Abstract
Can we teach Large Language Models (LLMs) to refrain from hallucinating factual statements? In this paper we present a fine-tuning strategy that we call ConfQA, which can reduce hallucination rate from 20-40% to under 5% across multiple factuality benchmarks. The core idea is simple: when the LLM answers a question correctly, it is trained to continue with the answer; otherwise, it is trained to admit "I am unsure". But there are two key factors that make the training highly effective. First, we introduce a dampening prompt—"answer only if you are confident"—to explicitly guide the behavior, without which hallucination remains high as 15%-25%. Second, we leverage simple factual statements, specifically attribute values from knowledge graphs, to help LLMs calibrate the confidence, resulting in robust generalization across domains and question types. Building on this insight, we propose the Dual Neural Knowledge framework, which seamlessly select between internally parameterized neural knowledge and externally recorded symbolic knowledge based on ConfQA’s confidence. The framework enables potential accuracy gains to beyond 95%, while reducing unnecessary external retrievals by over 30%.
[…]
Yin Huang at
1 Introduction
Despite the remarkable capabilities that Large Language Models (LLMs) have demonstrated, hallucination of factual statements remains a challenge (Maynez et al., 2020; Zhou et al., 2021; Ji et al., 2023a). It has been broadly realized that factual information shall not be fabricated or generalized, instead shall be anchored in internally parameterized neural knowledge or externally recorded symbolic content (stored in knowledge graphs, webpages, or other repositories). Significant progress has been made in both knowledge internalization through pre-training (Grattafiori et al., 2024) and external knowledge utilization via Retrieval-Augmented Generation (RAG) (Wei et al., 2021; Yu et al., 2022; Gao et al., 2024; Fan et al., 2024; Huang and Huang, 2024). However, a critical question remains: when should LLMs rely on parameterized knowledge versus external sources? In the absence of external data, we wish the LLM to state only reliably parameterized knowledge and eliminate hallucinations; in the presence of external data, we wish the LLM to conduct retrieval wisely to reduce latency, save resources, and avoid distractions (Jiang et al., 2023).
In this paper, we study this problem through three questions:
  • Q1. Does an LLM know what it knows?
  • Q2. Can we teach LLMs to refrain from hallucinations?
  • Q3. What is the optimal strategy for RAG triggering?
We conducted extensive experimental studies to answer these questions and made three contributions.
Figure 1: Overall factuality improvement of our Dual Neural Knowledge framework. Our fine-tuned model ConfQA reduces hallucination to under 5%; when combined with RAG, we increase accuracy by 45% on average while cutting latency by selective RAG triggering.
[…]
Our second contribution is an effective fine-tuning method for hallucination suppression, called ConfQA (Q2). Our training data checks the LLM’s inherent answer to a question, and teaches it to state “I am unsure” if the answer is incorrect. Though seemingly simple, there are two key ingredients to make the training effective. First, we discover an important dampener prompt—“Answer only if you are confident”, which plays a crucial role in guiding the LLM behavior. Second, the training data comprise exclusively of simple questions that inquire about an attribute of an entity; by concentrating the regularization on simple factual statements, which serves as building blocks to more complicated factual statements, the training teaches LLMs to generalize this behavior. ConfQA is able to reduce hallucinations to under 5% on various benchmarks (see Figure 1).
Our method exhibits strong transferability: although trained solely on questions derived from DBPedia, similar improvements are observed across other domains (e.g., IMDb), benchmarks (CRAG (Yang et al., 2024), SimpleQA (Wei et al., 2024a)), and long-form answers (LongFact (Wei et al., 2024b), Biography (Min et al., 2023)). Additionally, the fine-tuned model maintains its performance on general benchmarks such as MMLU (Hendrycks et al., 2021)(Section 5).
Our final contribution is an information-providing framework that we call the Dual Neural Knowledge (DualKnowl) framework (Q3). The framework simultaneously invokes both ConfQA and a RAG pipeline in parallel, but it halts the RAG pipeline early unless one of the two conditions is met: (1) when a question requests dynamic information—facts that may change over minutes, days, or years (Vu et al., 2023), or (2) when the ConfQA model responses with “I am unsure” to a factual question. Our empirical study shows that DualKnowl achieves QA accuracy comparable to always invoking RAG, while reducing latency by over 600ms on CRAG (Yang et al., 2024) (Section 6).
2 Related Work
Hallucination, with a systematic taxonomy introduced in Tonmoy et al. (2024), has been listed as a top open problem for Generative AI (Cao et al., 2023). There are two bodies of work on hallucination reduction: LLM post-training and RAG. We next discuss each in detail.
RAG with external knowledge:
Retrieval-augmented generation (RAG) (Lewis et al., 2020; Shi et al., 2024) has been extensively researched in academia and widely applied in industry to enhance LLMs’ capability to answer factual questions accurately; there have been numerous surveys on RAG, such as  Wei et al. (2021)Yu et al. (2022)Gao et al. (2024)Fan et al. (2024)Huang and Huang (2024). What is closely related to our work is RAG triggering strategies. Su et al. (2024) propose a strategy to trigger RAG when the generated token has high entropy or high self-reported attention. Similarly, Jiang et al. (2023) propose to use LLM internalized knowledge for generation, and apply RAG when the confidence of the next token is low.  Peng et al. (2023) proposes a system that conducts retrieval and answer generation iteratively, revising the LLM prompts to improve model responses using the factuality score from an automatic verifier. These methods all use token-level confidence, and often need multiple retrievals sequentially. Our work focuses on fact-level confidence, can apply in situations when token-level confidence is unavailable, and typically require a single retrieval.
LLM fine-tuning:
Training-based hallucination mitigation has been surveyed in Tonmoy et al. (2024). There are two directions for training: enriching the parameterized knowledge, and suppressing hallucinations.
A lot of study focuses on knowledge enrichment during post-training, such as using ground truths in QA data sets (Zhang et al., 2024), leveraging trustworthy sources and especially knowledge graphs (Bayat et al., 2023; Ji et al., 2023b), constructing factuality preference pairs for DPO (Rafailov et al., 2024) based on long form generation, and distilling from larger models (Elaraby et al., 2023). Recently it was discovered that feeding knowledge in post training would encourage the LLM to ignore what it learns during pre-training, thus leading to even more hallucinations (Lin et al., 2024; Gekhman et al., 2024). This is confirmed in our experiments.
On the other hand, a lot of research focuses fine-tuning on hallucination suppression. Sun et al. (2023b) teaches LLMs to recite factual passages to avoid hallucination. Dhuliawala et al. (2023) verifies responses with internalized knowledge before final generation. Tian et al. (2023) generates factuality preference rankings to favor factual statements consistent with external sources or internal knowledge. Xie et al. (2025) trains a factuality evaluator to provide LLM generators with claim-level factuality feedback. A few works focused on teaching LLMs about its confidence on factual statements, including pre-LLM work (Mielke et al., 2022; Kadavath et al., 2022) and recent work based on local intrinsic dimension (Yin et al., 2024).
The works that are closest to ours are Cheng et al. (2024) and Zhang et al. (2024), both instructing the LLM to refuse to answer questions where it has a low confidence about the answer. R-Tuning (Zhang et al., 2024) identifies such questions according to answer correctness and pads sure or unsure to the end of the answer. Cheng et al. (2024) in addition requires providing the correct answer consistently. Our ConfQA training is tremendously different in two ways: first, we use the dampener prompt, which reduces hallucination further by 5-11% in our empirical study; second, we focus on simple factual questions from the DBPedia knowledge graph, which increases factuality by up to 30%. Our experiments also show that requiring consistency in addition can cause a large correctness regression. We present the experimental comparison in detail in Section 5.
3 Experiment Setup
Benchmark Category Question types # Domain Size
Head-to-Tail (Sun et al., 2023a) short-form simple questions (attribute of entities) dbpedia, imdb 1,200
SimpleQA (Wei et al., 2024a) short-form general fact-seeking questions multiple domains 4,326
CRAG (Yang et al., 2024) short-form simple questions, reasoning questions 5 domains 642
LongFact (Wei et al., 2024b) long-form general questions 38 domains 250
AlpacaFact (Lin et al., 2024) long-form fack-seeking instruction-following multiple domains 241
Biography (Min et al., 2023) long-form biography questions celebrity 183
MMLU 5-shot (Hendrycks et al., 2021) general knowl. multi-choice questions 57 domains 14,042
MMLU pro (Wang et al., 2024) general knowl. multi-choice questions multiple domains 12,032
Table 1: The overall statistics of evaluation datasets.
Data sets:
As we focus on factuality, we wish to reduce hallucination on factual statements, without regressing performance on general knowledge and problem-solving tasks. We thus consider benchmarks in three categories: 1) Short-form factuality benchmarks, where the answers are mostly short; the question include both simple questions regarding an attribute of an entity, and complex ones that require comparison, aggregation, reasoning, and post-processing; 2) Long-form factuality benchmarks, where answers are expected to be long and contain multiple factual statements; 3) General knowledge benchmarks, which focuses on general knowledge and reasoning in diverse disciplines. Under these three categories we considered seven benchmarks, where Table 1 summarizes the benchmarks and Appendix 10 gives details.
Metrics and evaluation:
We use slightly different metrics to measure the factuality of answers generated by LLMs. For short-form factuality benchmarks, we compute the percentage of correct, incorrect, and missing (not attempted) answers, and then define two top-line metrics.
  • Factuality-score: Following the CRAG benchmark (Yang et al., 2024), we compute Factuality = correct% - incorrect%, which ranges from -1 to 1.
  • F1-score: Following the SimpleQA benchmark (Wei et al., 2024a), we compute F1-score as a harmonic mean of precision and recall, where Precision = correct% / (correct% + incorrect%), and Recall = correct%. F1-score ranges from 0 to 1.
Between the two metrics, F1-score is more lenient for incorrect answers (hallucinations), but factuality strongly prefers missing answers to hallucinations (see examples in Section 10). We decide the correctness of an answer by prompting Llama-3.1-405B to compare a generated answer with the ground truth. We use SimpleQA evaluation prompt 111 https://github.com/openai/simple-evals for the SimpleQA benchmark, and use Prompt 2 in Appendix 9 for other benchmarks. Similar to Yang et al. (2024), we observe a  99% accuracy in the judgment.
For long-form responses we use the automatic evaluation metric, VeriScore (Song et al., 2024), which computes precision, recall, and F1-score. The main difference from short-form questions is that we set the minimum number of facts required for perfect recall based on the median number of extracted claims per dataset, using their fine-tuned models for claim extraction and verification.
For general knowledge benchmarks, MMLU provides ground truths for the multi-choice questions. The score is computed as the percentage of correctly answered questions, as a weighted average among the 57 diverse subjects.
LLM Models and implementations:
We conduct experiments using six well-known LLMs: Llama3.1-8B, Llama3.1-70B (Touvron et al., 2023), GPT-4o-mini and GPT-4o (OpenAI, 2023), Claude3.5-Sonnet [^2] https://www.anthropic.com/news/claude-3-5-sonnet and Claude3.5-Haiku [^3] https://www.anthropic.com/claude/haiku. We conducted experiments on Nvidia H100 96GB HBM2e GPUs with different configurations based on different model sizes. For Llama3.1-70B models, we did fine tuning on 32 H100 GPUs and inference on 8 H100 GPUs.
4 Q1. Does an LLM Know What It Knows?
We first investigate whether an LLM can accurately estimate its own confidence in a factual statement. If so, we can rely on this estimation to choose between internalized neural knowledge and external symbolic knowledge. We measure confidence in two ways: self-reported confidence, and consistency of answers. We focus on the three short-form benchmarks: Head-to-Tail, SimpleQA, and CRAG. We investigate the calibration between confidence and QA accuracy; that is, when the model has a confidence of 0.8, is the QA accuracy close to 80%?
Figure 2: LLMs’ self-reported confidences is correlated with QA accuracy, but often over-confident. The answer consistency is often better calibrated with QA accuracy.
Confidence vs. Accuracy:
To obtain self-reported confidence, we prompt the LLM to directly provide a confidence score between 0 and 1 along with its answer (prompt template in Appendix 3). We remove missing answers, divide the reported confidences by equal-sized quantile bins, and plot the average accuracy within each bin. The top panels in Figure 2 show the calibration, leading to four observations. (We in addition compare calibration for head, torso, tail entities in Appendix 11).
  1. 1.
    The self-reported confidence is mostly positively correlated with QA accuracy, but LLMs tend to be over-confident (the correlation curves are below the ideal calibration dashed line). For example, when Llama-3.1-70B predicts a confidence of 80% on CRAG, the real accuracy is only 33%.
  2. 2.
    Notably, for the same model series, the smaller model is often more confident than the larger model (with an exception of Claude3.5 on CRAG), demonstrating the interesting correlation between ignorance and self-assurance.
  3. 3.
    Finally, the overconfidence is more pronounced when answering SimpleQA questions than on other benchmarks. A sample of 50 questions from SimpleQA shows that the questions are often nuanced for fairly popular entities (e.g. “What was the first line after the salutation in the letter sent to Sardar Patel by Abhay Charan De?", “In which month and year was Service Pack 3 for Windows Fundamentals for Legacy PCs released?”), possibly causing LLMs to be over-confident.
Consistency vs. Accuracy:
To measure consistency, we ask LLM the same question 20 times with the temperature set to 1.0, select the most frequent response as the final answer, and calculate its frequency among the 20 times as the consistency score. To be robust against minor differences, we determine the "most frequent" answer based on semantic similarity rather than exact string match.
The bottom panels in Figure 2 shows that consistency is mostly better calibrated than self-reported confidence, and largely overlays with the ideal calibration on CRAG. In addition, the calibration curve is more linear compared to self-reported confidence.
Summary:
Ideally, an LLM would know what it knows, give facts it has high confidence about, and refrain from including low-confidence facts to avoid hallucinations. Unfortunately, as this analysis shows, self-reported confidence tends to be over-confident, whereas answer consistency requires invoking LLMs multiple times and thus is impractical. We thus need to find a different strategy.
5 Q2. Can We Teach LLMs to Refrain from Hallucinating?
In this section, we shift gears and study whether we can fine tune an LLM to suppress hallucinations, by only answering a question that it has high confidence about, and admitting “I am unsure” otherwise. We have three key intuitions for this fine-tuning. First, the calibration study shows that LLM’s self-reported confidence is associated with answer accuracy but not calibrated, we thus teach it to calibrate its confidence by showing the ground truth. Second, we introduce a dampener prompt “Answer only if you are confident”, to explicitly guide LLM’s behavior. Third, as we wish to regularize behavior only for factual statements, we focus the teaching on atomic facts (attributes of entities) to avoid distractions of other factors.
5.1 Fine-tuning setup
Training:
We conducted supervised fine-tuning (SFT) and constructed the training data as a set of question-label pairs, where the label teaches the model the answer it shall generate.
The questions ask for atomic facts, and are generated from DBPedia, which covers a diverse set of domains (Intuition #3). We used the open-sourced script from Sun et al. (2023a) to generate question-answer pairs from DBPedia (exclusive from those in the Head-to-Tail benchmark), evenly distributed across head, torso, and tail entities.
We generated labels as follows. First, we prompted Llama-3.1-70B model to answer the DBPedia-based questions (Prompt 1 in Appendix 9). Then, we prompt Llama-3.1-405B to judge if the answer is consistent with the ground truth (Prompt 2 in Appendix 9). If the answer is correct, the label is the ground truth answer; otherwise, the label is “I am unsure about the answer” (Intuition #1).
Inferences:
We call our fine-tuned model ConfQA. At inference time, we prompt QA in two ways: without the dampener, and with the dampener, to examine the dampening effect. (Intuition #2).
Other solutions:
We compare our solution with two state-of-the-art solutions. R-Tuning (Zhang et al., 2024) generates its training data by adding a prompt “Are you sure you accurately answered the question based on your internal knowledge?” in the question, and padding “I am sure” or “I am unsure” based on correctness of the generated answer. In the inference, it again appends the prompt and applies post-processing by removing answers with the suffix of “I am unsure”. We used both MMLU (proposed in the paper) and DBPedia for training. IDK (Cheng et al., 2024) requires answer consistency in addition to answer correctness, and we add a consistency requirement of at least four out of five times. We used DBPedia for its fine-tuning for more direct comparison.
Additionally, we considered the following alternatives for ablation study; more alternatives are discussed in Appendix 14.
Implementation and hyper-parameters:
We experimented with Llama-3.1-70B (we observed similar trend on Llama-3.1-8B). We conducted a simple scaling-law study (see Appendix 12) and decided to run one epoch on 3K high quality training data, with a learning rate of 1e-6 and a batch size of 1.
5.2 Results on short-form question answers
Model Corr (Rec) Miss Hall. Fac. Prec F1 Corr (Rec) Miss Hall. Fact. Prec F1
DBpedia (in-domain) IMDB (out-of-domain)
Llama-3.1 52.0 22.0 26.0 26.0 66.7 58.4 44.8 34.2 21.0 23.8 68.1 54.1
Llama-3.1 (D) 47.0 26.8 26.2 20.8 64.2 54.3 40.7 36.2 23.2 17.5 63.7 49.6
R-tuning (MMLU) 24.3 67.3 8.3 16.0 74.5 36.6 28.2 60.5 11.3 16.9 71.3 40.4
R-tuning (DBPedia) 24.5 67.8 7.7 16.8 76.2 37.1 25.3 70.2 4.5 20.8 84.9 39.0
IDK (DBPedia) 17.0 81.5 1.5 15.5 91.9 28.7 22.0 77.0 1.0 21.0 95.7 35.8
ConfQA 49.9 33.5 17.5 31.5 74.0 59.0 43.0 42.0 16.0 27.0 73.0 54.0
ConfQA (D) 31.5 63.3 5.2 26.3 85.9 46.1 32.5 63.3 4.2 28.3 88.6 47.6
SimpleQA (out-of-domain) CRAG (out-of-domain)
Llama3.1 20.0 44.1 35.9 -15.8 35.8 25.7 58.7 15.6 25.7 33.0 69.6 63.7
Llama3.1 (D) 16.8 48.0 35.2 -18.4 32.3 22.1 57.5 22.3 20.2 37.2 73.9 64.7
R-tuning (MMLU) 20.3 38.0 41.7 -21.4 32.8 25.1 57.8 17.1 25.1 32.7 69.7 63.2
R-tuning (DBPedia) 3.7 83.3 13.0 -9.3 22.0 6.3 31.6 55.0 13.4 18.2 70.2 43.6
IDK (DBPedia) 0.6 99.1 0.2 0.4 73.0 1.2 20.7 78.2 1.1 19.6 95.0 34.0
ConfQA 17.3 55.8 26.8 -9.5 39.2 24.0 57.0 19.6 23.4 33.6 70.9 63.0
ConfQA (D) 4.9 93.1 2.1 2.8 70.2 9.1 39.4 56.2 4.4 35.0 90.0 54.8
Table 2: Overall factuality improvement on short-form benchmarks; ConfQA can reduce hallucination to below 5% with the dampener prompt. All numbers are in percentage (%).
Overall results:
Table 2 presents our overall results, and shows interesting effect of the dampener prompt. First, without fine-tuning, the effect of the dampener is inconsistent. For all benchmarks, the dampener increased the percentage of missing answers by 2-7%. However, it (correctly) dampens hallucinations on CRAG, but dampens correct answers and thus reduced the factuality on Head-to-tail and SimpleQA. This is not surprising since the LLM confidence is not well calibrated.
Second, after the fine-tuning, ConfQA improves precision by up to 40% and improves factuality by up to 20%. Without the dampener, we observe mild reduction of the hallucinations, accompanied with minor sacrifice on correct answers, leading to factuality improvement up to 6%. With the dampener, the LLM is much more conservative and the hallucination rate drops to below 5% on all benchmarks. As a side effect, correctness also drops significantly; for example, since SimpleQA focuses on nuanced facts, after the finetuning we observe nearly zero correctness. Still, the factuality increases by up to 18% compared to the baseline.
Third, R-tuning mostly has higher precision and lower hallucination, especially if trained on DBPedia. However, we also observe much lower correctness. We suspect this is because when the model gives a wrong answer, it feeds ground truths as additional knowledge and causes confusion; as we will discuss soon, we observe similar trend for Fact-feeding, with a similar flavor. We also observe better results when trained on DBPedia than on MMLU, which mixes factual statements and skills and can cause confusion.
Fourth, IDK obtains the lowest hallucination rate (below 1.5% for all benchmarks), as it requires in addition the consistency signal and thus is stricter. However, the correctness also drops significantly, reducing overall factuality.
Fifth, even though the training data are generated only from DBPedia, ConfQA behavior changes on the other datasets as well, showing amazing generalization.
Finally, we note that although the factuality metric increases most of the time, the F1-metric stays flat or drops, as the latter is more lenient to hallucinations as discussed in Appendix 10.
Model DBpedia (in-domain) IMDB (out-of-domain) SimpleQA (out-of-domain) CRAG (out-of-domain)
Corr Miss Hall Fac Corr Miss Hall Fac Corr Miss Hall Fac Corr Miss Hall Fac
Llama-3.1 47.0 26.8 26.2 20.8 40.7 36.2 23.2 17.5 16.8 48.0 35.2 -18.4 57.5 22.3 20.2 37.2
ConfQA 31.5 63.3 5.2 26.3 32.5 63.3 4.2 28.3 4.9 93.1 2.1 2.8 39.4 56.2 4.4 35.0
* No-dampener 36.0 50.2 13.8 22.2 34.2 50.7 15.2 19.0 5.8 87.5 6.7 -0.9 46.0 44.2 9.8 36.2
* MMLU-as-source 8.2 89.8 2.0 6.2 16.7 82.0 1.3 15.4 0.6 98.8 0.5 0.1 7.0 92.7 0.3 6.7
* GT-as-label 48.0 2.8 49.2 -1.2 41.2 4.3 54.5 -13.3 17.8 13.7 68.4 -50.6 53.7 14.2 32.1 21.6
* Fact-feeding 20.7 76.7 2.7 18.0 25.5 70.7 3.8 21.7 2.5 94.7 2.8 -0.3 22.4 74.5 3.1 19.3
Table 3: Ablation study when applying dampening in inference, showing effectiveness of our fine-tuned model. All numbers are in percentage (%) and full results in Table 8.
Ablation study
We now compare ConfQA with the several alternatives, as shown in Table 3 (full comparison in Table 8 in Appendix 14). MMLU-as-source obtains low hallucination (below 2%) but also significantly lower correctness. We suspect this is because MMLU contains a diverse set of tasks, reducing overall confidence of the model.
On the other hand, No-dampener increases correctness and reduces missing rate, but can increase hallucinations, showing the important role of the dampener in training as well. GT-as-label achieves the highest correctness and lowest missing rate, but becomes over-confident to hallucinate (hallucination rate can reach 70%). This is consistent with observations in previous work (Lin et al., 2024; Gekhman et al., 2024) that feeding facts in the SFT-stage can teach LLMs to hallucinate.
Interestingly, Fact-feeding combines our ConfQA fine-tuning data with Tulu facts, drops hallucinations but also correctness, similar to R-tuning. We suspect this is because what our training data teach the LLM (saying unsure) is of different purpose from what the extra Tulu facts teach the LLM (feeding knowledge), when mixed together can offset each other and cause confusion.
(a) DBPedia
(b) IMDB
(c) CRAG
Figure 3: ConfQA suppresses more on long-tail facts than popular facts.
Answer distributions:
Finally, we show in Figure 3 the distribution of correct, missing, and incorrect answers for entities of different popularity, before and after fine-tuning, with and without dampening. It confirms that fine-tuning suppresses hallucinations, and the dampener prompt further reduces hallucinations. Additionally, it shows ConfQA suppresses more on long-tail facts, where it lacks confidence.
Model Long Fact Alpaca Fact Biography
Prec Rec F1 Miss Prec Rec F1 Miss Prec Rec F1 Miss
Llama3.1 64.5 65.4 64.3 0 62.3 71.0 63.8 0 35.4 40.3 37.1 0
RAG (Llama3.1) (Yu et al., 2022) 71.7 74.6 72.7 0 65.8 74.3 66.0 0 44.9 48.1 43.8 0
ConfQA 67.0 67.7 66.7 0.8 62.2 71.1 63.8 0.4 42.0 46.5 42.6 12.6
Table 4: ConfQA improves precision and recall for long-form answer generation.
5.3 Results on other benchmarks
On the long-form benchmarks, we do not apply the dampener prompt, and instead we retrieve 10 passages using the input prompts as queries and append “Provide as many specific details and examples as possible (such as names of people, numbers, events, locations, dates, times, etc.)” to the end of the prompt to encourage the model to provide as much confident information as possible. Table 4 shows that ConfQA achieves higher or comparable precision and recall, except for 13% biography questions where it has low confidence about and does not answer. For RAG, we use Contriever (Izacard et al., 2022) to retrieve passages from C4 (Raffel et al., 2020) and Wikipedia, following the setting in MassiveDS (Shao et al., 2024).
Model MMLU (5-shot) MMLU-Pro
Llama3.1 82.7 66.3
ConfQA 82.8 65.4
Table 5: ConfQA does not regress on MMLU.
Figure 4: RAG invocation architecture.
We also evaluate ConfQA on the standard MMLU benchmark. Table 5 shows that the scores are mostly similar to the baseline.
Summary:
Our experiments validated the effectiveness of our ConfQA fine-tuning strategy. If one aims to eliminate hallucinations, we recommend applying ConfQA with the dampener; if one aims to maintain the correctness and meanwhile reduce hallucinations as much as possible, we recommend ConfQA without the dampener.
6 Q3. What Is the Optimal Strategy for Triggering RAG?
Since our fine-tuned model can reduce hallucinations to nearly zero, we can invoke the RAG pipeline when it says unsure, and rely on the LLM-generated answer otherwise. Figure 4 depicts a RAG invocation architecture. The system always answers a dynamic question that inquires changing information through the RAG pipeline. When a static question arrives, the system invokes LLM generation and the RAG pipeline in parallel. When LLM generates a non-missing answer, the system early-stops the RAG pipeline and outputs the answer; otherwise, the system waits and outputs the answer from the RAG pipeline.
Calculations based on Table 2 shows that our framework enables potential accuracy gains to beyond 95%, while reducing unnecessary external retrievals by over 30%. We next evaluate QA quality and latency through a real RAG implementation, which invokes search APIs (Bing API and Knowledge Graph API) for retrieval, and passes the retrieved content to Llama-3.1-70B to generate the responses. Table 6 reports the end-to-end QA accuracy and latency for our proposed RAG architecture, and compares it with not invoking RAG and invoking RAG everywhere. ConfQA-triggered RAG obtains similar quality to RAG-everywhere, but cut latency by 300ms P50 and 400ms P90 for CRAG. The latency-cut is less pronounced on SimpleQA since RAG should be triggered for the majority of the questions. The quality improvement is small on CRAG because it contains a lot of complex questions requiring reasoning over retrieval results, which our simple RAG implementation does not excel.
Model SimpleQA CRAG
Upper Corr Miss Hall Fac L-P50 L-P90 Upper Corr Miss Hall Fac L-P50 L-P90
LLM-only 20.0 20.0 44.1 35.9 -15.8 480 896 58.7 58.7 15.6 25.7 33.0 480 896
RAG-everywhere 100.0 78.1 11.5 10.5 67.6 1,900 2,780 100.0 61.1 15.1 23.8 37.3 1,900 2,780
ConfQA-based 95.1 77.2 11.4 11.5 65.7 1,802 2,650 95.6 62.3 14.2 23.5 38.8 1,278 1,955
Table 6: ConfQA-based RAG invocation achieves similar quality to RAG-everywhere, while cutting latency.
7 Limitations
Our experiments focus on SFT, and we leave DPO-based fine-tuning for future work. We conducted training on the Llama-3.1 model of different sizes; it would be interesting to experiment with other open sourced models and later versions of models. We empirically compared DBPedia and MMLU, where the former contains only simple factual questions, and the latter contains questions ranging from factual to reasoning. A comprehensive study regarding the effect of sources with different coverage in this spectrum would improve the understanding. We can extend our learnings beyond factual statements, for math, coding, reasoning etc. Lastly, SFT requires the access to LLM itself for fine tuning, and thus restrict the application of the proposed framework to proprietary LLMs only accessible through APIs.
8 Conclusion
Recent studies have shown that LLMs acquire substantial knowledge during pre-training, and that introducing new knowledge during post-training can often increase hallucinations (Lin et al., 2024; Gekhman et al., 2024). Rather than injecting additional knowledge, the Dual Neural Knowledge (DualKnowl) framework we propose trains LLMs to identify and withhold low-confidence factual claims, deferring to RAG in such cases to improve factuality. Through a comprehensive set of experiments, we show that 1) LLMs tend to be over-confident on what they know; 2) we can fine-tune LLM to refrain from generating inconfident factual statements and thus reducing the hallucination rate to below 5%; and 3) using this fine-tuned model for RAG-triggering can reach similar answer accuracy as RAG-everywhere, whereas reducing RAG retrievals by over 30% to save cost. Our proposed framework naturally integrates internal neural knowledge with external symbolic knowledge, allowing for improvements in both factuality and latency as LLMs develop richer internal knowledge and more effective RAG capabilities.